Frank has decided to invest $8000 into two different funds. His RRSP pays interest at a rate of 9%/a, and his GASP pays at a rate of 6%/a. If he earned $615 of interest after one year, how much did he originally invest into each account?


1
Expert's answer
2021-05-21T11:09:00-0400

Say, x amount is invested at 9% interest rate.

Say the interest earned = I1

"I1 = \frac{x * 9}{100}"

"Or I1 = \frac{9x}{100}"


And, Therefore, 8000-x will be invested at 6% interest rate.

Say the interest earned = I2

"I2 = \frac{(8000-x) * 6}{100}"

"Or I2 = \frac{48000 - 6x)}{100}"


Given, I1 + I2 = 615


"\frac{9x}{100} + \frac{48000-6x}{100} = 615"


Or

9x + 4800 - 6x = 61500

Or

3x = 61500 - 48000

Or

x = 4500


Final Answer: Amount x = 4500 is invested at 9% and balance 3500 ( = 8000-4500) is invested at 6%.